OpenStudy (anonymous):

7. A local weather forecaster is accurate 85% of the time when predicting precipitation for the day. What is the probability that she will make correct precipitation predictions 4 days in a row? Round to the nearest whole percent. (2 points)

Parth (parthkohli):

If they make it 85% correctly, then to pick up with four days, they'll just do it 85% accurately each day. Well this makes sense. So the probability would be 85% multiplied to itself four times. \(\Large \color{Black}{\Rightarrow 0.85 \times 0.85 \times 0.85 \times 0.85}\)

OpenStudy (anonymous):

8. A bag contains 6 purple marbles and 7 white marbles. Two marbles are drawn at random. One marble is drawn and not replaced. Then a second marble is drawn. What is the probability that the first marble is white and the second one is purple? (2 points)